OpenStudy (anonymous):

Part A: Explain why x = 6 makes 4x − 5 ≤ 19 true but not 4x − 5 < 19. Part B: What value from the set {4, 5, 6, 7, 8} makes the equation 4x + 3 = 23 true? Show your work.

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

automaticloveletter

OpenStudy (anonymous):

@automaticloveletter

OpenStudy (anonymous):

@Phil1111

OpenStudy (anonymous):

@cookielate

OpenStudy (anonymous):

Help

OpenStudy (automaticloveletter):

Don't get A. Part B: What value from the set {4, 5, 6, 7, 8} makes the equation 4x + 3 = 23 true? Show your work. 4*4+3=19 x 5*4+3=23 (correct) 6*4+3=27 x 7*4+3=31 x 8*4+3=25 x

OpenStudy (automaticloveletter):

Part A: Explain why x = 6 makes 4x − 5 ≤ 19 true but not 4x − 5 < 19. 4 * 6 = 24 - 5 = 19 and 19 is equal to 19, that's why it's true. 4 * 6 - 5 = 19 and 19 is not less than 19 that's why it's false.

OpenStudy (automaticloveletter):

Hope I helped :)

OpenStudy (anonymous):

yes thank you